Find the solution to the given system that satisfies the given initial condition.Find the solution to the given system that satisfies the given initial condition. -7 -1 x(t) = x(t), 2-5 - 5 2 (a) (0) = (b)
Answer #1

x(t) = -7 -L X (+) be the given system of differential 2. -5 equation, or X(+) = A x(+) Now characteristic equation of 1817X(t) = (6+2)t e 79 est et [1 [ - i=-1 it [cost ti sen(t)] [121] .66 Cosct) sin(t) ti(cos(t) - since) 2 cos ct) i zsen(t) 792Now put initial condition XCO) = 13 then [8]= [2]+ e[:] coso=t, Sinozo, e=1 - 5. X0= Is] is implies • 2¢0 7 F0 and e thereforX(t) 6t la cos(t) = sin(t) 169]-* =*[aunch Cos (t)-sin(t) 2 COSCL) X(-21) a pencatat [1] 1 - Setari then from Go. cos €217) -d X(72) = then from -6:57 & Ge [] (i Egers [cos27%) -im (90 ] Töltse cos) sin (2) 2 sin (2) como 2 Coscry) 317 qe -30 [2] Cos